For the problem f(x)=x-3 and g(x)=x^2 +1
A) what is (f∘g)(-1)?
B) what is (g∘f)(x)?

Answer by stanbon(75887)

For the problem f(x)=x-3 and g(x)=x^2 +1
A) what is (f∘g)(-1)?
fog(x) = f[g(x)] = f[x^2+1] = x^2+1-3 = x^2-2
----------------------------------------
B) what is (g∘f)(x)?
gof(x) = g[f(x] = g[x-3] = (x-3)^2+1 = x^2-6x+9+1 = x^2-6x+10
-----------
Cheers,
Stan H.
------------